A container shaped as a rectangular prism has a volume of 124 cubic feet. If the length is doubled but the other dimensions remain the same, what will be the volume of the new cube?
A. 248 cu ft
B. 372 cu ft
C. 496 cu ft
D. 992 cu ft
What would the formula be? Would you double the length? How do you find this?

Answer by fcabanski(1391)

The volume of a rectangular prism (a rectangular box) is Length x Width x Height = L*W*H


Double the Length is 2L. So the new volume is 2L*W*H which is double the original volume.


Let's use a simple example of a rectangular prism with L = 2, W=3 and H=1


V = 2*3*1 = 6. If the length is doubled to be 4, the volume is 4*3*1 = 12. It doubles, because a factor of 2 was added to only one of the dimensions.


The new volume is 2L*W*H = 2*Old Volume = 248 cubic feet.